# RL3: Boosting Meta Reinforcement Learning via RL inside RL2

### v4

@@ 7 Conclusion
In this paper, we introduced RL3, a principled hybrid approach that combines the strengths of traditional RL and meta-RL and provides a more robust and adaptable reinforcement learning algorithm for complex and diverse environments. We advanced intuitive and theoretical arguments regarding its suitability for meta-RL and presented empirical evidence to validate those ideas. Specifically, we demonstrated that RL3 outperforms RL2 across a wide range of tasks, enhancing long-term performance and generalization on out-of-distribution tasks, while maintaining efficiency in the short term. In future work, we plan to explore extending RL3 to handle continuous state spaces, possibly using state abstractions for discretization.
@@ A.3 Object-level Q-estimates and Meta-level Values
Proof of Equation (4): In standard meta-RL, the only observed variable in the POMDP state $\bar{s}_{t}=[s_{t},i]$ at time $t$ is the state $s_{t}$ of the current MDP i.e., $\bar{\omega}_{t}=s_{t}$, while the task identity $i$ is hidden. However, in RL3, $\bar{\omega}_{t}$ includes the vector of $Q$-estimates $Q^{t}_{i}(s_{t})$ for the hidden task, which means that the meta-level observation function $\bar{O}(\bar{\omega}|\bar{b},a)$ factors in the probability that a particular $Q$-esimate will be observed following an action $a$ given an initial belief $\bar{b}$ state. (Note that we will use $\bar{b}(\bar{s})$ and $\bar{b}(i)$ interchangeably since $i$ is the only hidden variable in $\bar{s}$). In practice, such $Q$-value estimates provide excellent evidence (see Appendix D) for task identification. This allows for robust belief recovery even if the initial belief is not Bayes-optimal (or altogether not maintained), especially as the Q-estimates converge and stabilize in the limit, leading to two cases:
@@
This proves equation (4). Note that in the limit, the task can be identified perfectly from the stream of experiences as all state-action pairs are explored, and the meta-level value function becomes equivalent to the optimal object-level value function of the identified (or current) task. However, the above proof demonstrates that RL3 can infer this equivalency implicitly in the limit without relying on the stream of experiences or identifying the task fully, and furthermore, directly model the meta-value function in terms of the supplied object-level value function. ∎
Proof of Equation (6): We first write the Bellman equation for the optimal meta-level POMDP value function in its belief-MDP representation:
@@ B.1 RL2
Our modified implementation of RL2 uses transformer decoders Vaswani et al. [2017] instead of RNNs to map trajectories to action probabilities and meta-values, in the actor and the critic, respectively, and uses PPO instead of TRPO for outer RL. The decoder architecture is similar to Vaswani et al. [2017], with 2 layers of masked multi-headed attention. However, we use learned position embeddings instead of sinusoidal, followed by layer normalization. Our overall setup is similar to Esslinger et al. [2022].
For each meta-episode of interactions with an MDP $M_{i}$, the actor and the critic transformers look at the entire history of experiences up to time $t$ and output the corresponding action probabilities $\pi_{1}...\pi_{t}$ and meta-values $\bar{V}_{1}...\bar{V}_{t}$, respectively. An experience input to the transformer at time $t$ consists of the previous action $a_{t-1}$, the latest reward $r_{t-1}$, the current state $s_{t}$, episode time step $t_{\tau}$, and the meta-episode time step $t$, all of which are normalized to be in the range $[0,1]$. In order to reduce inference complexity, say at time step $t$, we append $t$ new attention scores (corresponding to experience input $t$ w.r.t. the previous $t-1$ experience inputs) to a previously cached $(t-1)\times(t-1)$ attention matrix, instead of recomputing the entire $t\times t$ attention matrix. This caching mechanism is implemented for each attention head and reduces the inference complexity at time $t$ from $\mathcal{O}(t^{2})$ to $\mathcal{O}(t)$. Note that this caching mechanism is possible only when the input to the transformer is always the entire history, built up incrementally, instead of a moving window, since removing early experiences from the moving window would invalidate the cached attention scores as they attend to the entire history.
[A note on using a moving context to apply a model trained $H=500$ to $H=1000$: A naive implementation would use a moving window context of 500 time steps. However, with this choice, a model highly biased towards exploiting rather than exploring near the end of the window would only exploit from time steps 501 through 1000 even though there is room more exploration. As a workaround to this problem, we use a context length of 250 steps until time step 750, and increase it gradually thereafter to 500 steps towards the end. In our experiments, this heuristic rule led to a significant improvement over the naive method.]
Note that our results are not significantly better than those in the original RL2 paper, which uses a recurrent neural network, for the Bandits and the MDPs domain. However, our transformer version performs better on the GridWorld domain. Moreover, the transformer version drastically reduces real-world training time and is more stable and easier to tune.
@@ B.2 RL3
The transformer input in RL3 includes of a vector of $Q$-estimates and a vector of action counts at each step $t$ for the corresponding state. As mentioned in Section 4.2, this is implemented in our code by simply converting MDPs in the problem set to VAMDPs using a wrapper, and running our implementation of RL2 thereafter. The Markov version of RL3 uses a dense neural network, with two hidden layers of 64 nodes each, with ReLU activation function.
For object-level RL, we use model estimation followed by value iteration (with discount factor $\gamma=1$) to obtain $Q$-estimates. The transition probabilities and the mean rewards are estimated using maximum likelihood estimation (MLE). Rewards for unseen actions in a given state are assumed to be 0. States are added to the model incrementally when they are visited, so that value iteration does not compute values for unvisited states. Moreover, value iteration is carried out only for iterations equal to the task horizon (which is 1, 10, 250, 350 for Bandits, MDPs, 11x11 GridWorld, 13x13 GridWorld domains, respectively), unless the maximum Bellman error drops below 0.01.
@@ B.3 RL3-coarse
Table 5: RL2 /RL3 Hyperparameters
Hyperparameter Value
Learning Rate (Actor and Critic) 0.0003 (Bandits, MDPs)
0.0002 (GridWorlds)
Adam $\beta 1,\beta 2,\epsilon$ $0.9,0.999,10^{-7}$
Weight Decay (Critic Only) $10^{-2}$
Batch size 32000
Rollout Length Interaction Budget ($H$)
Number of Parallel Envs Batch Size $\div$ $H$
Minibatch Size 3200
Entropy Regularization Coeff 0.001 (Bandits $H=100$)
0.04 (GridWorlds)
0.01 otherwise
PPO Iterations See training curves
Epochs Per Iteration 8
Max KL Per Iteration 0.01
PPO Clip $\epsilon$ 0.2
GAE $\lambda$ 0.3
Discount Factor $\gamma$ 0.99
Decoder Layers 2
Attention Heads 4
Activation Function gelu
Decoder Size ($d\_{model}$) 64
Figure 4: Average meta-episode return vs PPO iterations for Bandits $H=100$ and $H=500$
Figure 5: Average meta-episode return vs PPO iterations for MDPs $H=100$ (left) and $H=500$ (right)
@@
C Training Curves
Figs. 4, 5, and 6 show the training curves for Bandits, MDPs, and GridWorld environments, respectively, across 4 random seeds. The results in the main text may differ slightly since the actor models were evaluated greedily, whereas the training curves reflect the actors’ stochastic policies.
We ran each experiment on an NVIDIA GeForce RTX 2080 Ti GPU, which took approximately 12-24 hours for Bandits and MDPs ($H=100$); and took approximately 3-4 days for other domains.
@@ E.1 Bernoulli Multi-Armed Bandits
We use the same setup described in Duan et al. [2016]. At the beginning of each meta-episode, the success probability corresponding to each arm is sampled from a uniform distribution $\mathcal{U}(0,1)$. To test OOD generalization, we sample success probabilities from $\mathcal{N}(0.5,0.5)$
@@ E.2 Random MDPs
We use the same setup described in Duan et al. [2016]. The MDPs have 10 states and 5 actions. For each meta-episode, the mean rewards $R(s,a)$ and transition probabilities $T(s,a,s^{\prime})$ are initialized from a normal distribution ($\mathcal{N}(1,1)$) and a flat Dirichlet distribution ($\alpha=1$), respectively. Moreover, when an action $a$ is performed in state $s$, a reward is sampled from $\mathcal{N}(R(s,a),1)$. To test OOD generalization, the reward function is made deterministic and initialized from $\mathcal{U}(0,2)$

### v5

@@ 7 Limitations and Conclusion
To conclude, in this paper, we introduced RL3, a principled hybrid approach that combines the strengths of traditional RL and meta-RL and provides a more robust and efficient meta-RL algorithm. We advanced intuitive and theoretical arguments regarding its suitability for meta-RL and presented empirical evidence to validate those ideas. Specifically, we demonstrated that RL3 holds potential to enhance long-term performance, generalization on out-of-distribution tasks and reducing meta-training time. In future work, we plan to explore extending RL3 to handle continuous state spaces.
@@ A.3 Object-level Q-estimates and Meta-level Values
Proof of Equation 4: In standard meta-RL, the only observed variable in the POMDP state $\bar{s}_{t}=[s_{t},i]$ at time $t$ is the state $s_{t}$ of the current MDP i.e., $\bar{\omega}_{t}=s_{t}$, while the task identity $i$ is hidden. However, in RL3, $\bar{\omega}_{t}$ includes the vector of $Q$-estimates $Q^{t}_{i}(s_{t})$ for the hidden task, which means that the meta-level observation function $\bar{O}(\bar{\omega}|\bar{b},a)$ factors in the probability that a particular $Q$-esimate will be observed following an action $a$ given an initial belief $\bar{b}$ state. (Note that we will use $\bar{b}(\bar{s})$ and $\bar{b}(i)$ interchangeably since $i$ is the only hidden variable in $\bar{s}$). In practice, such $Q$-value estimates provide excellent evidence (see Appendix D) for task identification. This allows for robust belief recovery even if the initial belief is not Bayes-optimal (or altogether not maintained), especially as the Q-estimates converge and stabilize in the limit, leading to two cases:
@@
This proves equation 4. Note that in the limit, the task can be identified perfectly from the stream of experiences as all state-action pairs are explored, and the meta-level value function becomes equivalent to the optimal object-level value function of the identified (or current) task. However, the above proof demonstrates that RL3 can infer this equivalency implicitly in the limit without relying on the stream of experiences or identifying the task fully, and furthermore, directly model the meta-value function in terms of the supplied object-level value function. ∎
Proof of Equation 6: We first write the Bellman equation for the optimal meta-level POMDP value function in its belief-MDP representation:
@@ B.1 RL2
Our modified implementation of RL2 uses transformer decoders [37] instead of RNNs to map trajectories to action probabilities and meta-values, in the actor and the critic, respectively, and uses PPO instead of TRPO for outer RL. The decoder architecture is similar to [37], with 2 layers of masked multi-headed attention. However, we use learned position embeddings instead of sinusoidal, followed by layer normalization. Our overall setup is similar to [55].
For each meta-episode of interactions with an MDP $M_{i}$, the actor and the critic transformers look at the entire history of experiences up to time $t$ and output the corresponding action probabilities $\pi_{1}...\pi_{t}$ and meta-values $\bar{V}_{1}...\bar{V}_{t}$, respectively. An experience input to the transformer at time $t$ consists of the previous action $a_{t-1}$, the latest reward $r_{t-1}$, the current state $s_{t}$, episode time step $t_{\tau}$, and the meta-episode time step $t$, all of which are normalized to be in the range $[0,1]$. In order to reduce inference complexity, say at time step $t$, we append $t$ new attention scores (corresponding to experience input $t$ w.r.t. the previous $t-1$ experience inputs) to a previously cached $(t-1)\times(t-1)$ attention matrix, instead of recomputing the entire $t\times t$ attention matrix. This caching mechanism is implemented for each attention head and reduces the inference complexity at time $t$ from $\mathcal{O}(t^{2})$ to $\mathcal{O}(t)$.
@@ B.2 RL3
The input of the transformer in RL3 includes a vector of $Q$ estimates (in practice, they are supplied as the vector of advantage estimates ($Q-\text{max}_{a}Q$) along with the value function ($\text{max}_{a}Q$) separately) and a vector of action counts at each step $t$ for the corresponding state. As mentioned in Section 4.2, this is implemented in our code simply by converting MDPs in the problem set to VAMDPs using a wrapper and running our implementation of RL2 thereafter. The pseudocode is shown in the algorithm 1. The Markov version of RL3 uses a dense neural network, with two hidden layers of 64 nodes each, with the ReLU activation function.
Algorithm 1 Value-Augmenting Wrapper for Discrete MDPs
procedure ResetMDP(vamdp)
    vamdp.$t\leftarrow 0$; vamdp.$t_{\tau}\leftarrow 0$
    vamdp.rl $\leftarrow\textsc{InitRL}()$
    $s=\textsc{ResetMDP}$(vamdp.mdp)
    return $\textsc{OneHot}(s)\cdot Q[s]\cdot N[s]$
procedure StepMDP(vamdp, $a$)
    $s\leftarrow$ mdp.$s$
    $r,s^{\prime}\leftarrow\textsc{StepMDP}$(vamdp.mdp, $a$)
    $d\leftarrow\textsc{Terminated}$(vamdp.mdp)
[…]
    vamdp.$Q\leftarrow\textsc{UpdateRL}$(vamdp.rl, $s,a,r,s^{\prime},d$)
[…]
         vamdp.$t_{\tau}\leftarrow 0$
         $s^{\prime}\leftarrow\textsc{ResetMDP}$(vamdp.mdp)
    return $r,$ $\textsc{OneHot}(s^{\prime})\cdot Q[s^{\prime}]\cdot N[s^{\prime}]$triangleright Concatenate state, $Q$-estimates and action counts
procedure Terminated(vamdp)
    return vamdp.$t\geq H$
For object-level RL, we use model estimation followed by value iteration (with discount factor $\gamma=1$) to obtain $Q$-estimates. The transition probabilities and the mean rewards are estimated using maximum likelihood estimation (MLE), with Laplace smoothing (coefficient = 0.1) for transition probabilities estimation. For unseen actions, rewards are assumed to be zero, and transitions equally likely to other states. States are added to the model incrementally when they are visited, so that value iteration does not compute values for unvisited states. Moreover, value iteration is carried out only for iterations equal to the task horizon (which is 1, 10, 250, 350 for Bandits, MDPs, 11x11 GridWorld, 13x13 GridWorld domains, respectively), unless the maximum Bellman error drops below 0.01.
@@ B.3 RL3-coarse
(a) $H=128$
(b) $H=256$
(c) $H=512$
(d) $H=1024$
(e) $H=1500$
(f) $H=2048$
Figure 5: Average meta-episode return vs PPO iterations for MDPs domain for different interaction budgets.
@@
C Training
Figs. 5, and 6 show the training curves for MDPs, and GridWorld environments, respectively, across 3 random seeds. The results in the main text correspond to the median model. We ran the experiments on Nvidia GeForce RTX 2080 Ti GPUs for context length $\leq 256$ which took approximately 12-24 hours, and on Nvidia A100 GPUs for higher context lengths, which took 1-2 days.
@@ E.1 Bernoulli Multi-Armed Bandits
We use the same setup described by Duan et al. [4]. At the beginning of each meta-episode, the success probability corresponding to each arm is sampled from a uniform distribution $\mathcal{U}(0,1)$. To test OOD generalization, we sample success probabilities from $\mathcal{N}(0.5,0.5)$
@@ E.2 Random MDPs
We use the same setup described by Duan et al. [4]. The MDPs have 10 states and 5 actions. For each meta-episode, the mean rewards $R(s,a)$ and transition probabilities $T(s,a,s^{\prime})$ are initialized from a normal distribution ($\mathcal{N}(1,1)$) and a flat Dirichlet distribution ($\alpha=1$), respectively. Moreover, when an action $a$ is performed in state $s$, a reward is sampled from $\mathcal{N}(R(s,a),1)$. To test OOD generalization, the transition probabilities are initialized with Dirichlet $\alpha=0.25$.
@@ E.3 GridWorlds
Table 3: RL2 /RL3 Hyperparameters
Hyperparameter Value
Learning Rate (Actor and Critic) 0.0003 (Bandits, MDPs)
0.0002 (GridWorlds)
Adam $\beta 1,\beta 2,\epsilon$ $0.9,0.999,10^{-7}$
Weight Decay (Critic Only) $10^{-2}$
Batch size 32768
Rollout Length Interaction Budget ($H$)
Number of Parallel Envs Batch Size $\div$ $H$
Minibatch Size 4096
Entropy Regularization Coeff 0.1 with decay (MDPs)
0.04 (GridWorlds)
0.01 (Bandits)
PPO Iterations See training curves
Epochs Per Iteration 8
Max KL Per Iteration 0.01
PPO Clip $\epsilon$ 0.2
GAE $\lambda$ 0.3
Discount Factor $\gamma$ 0.99
Decoder Layers 2
Attention Heads 4
Activation Function gelu
Decoder Size ($d\_{model}$) 64